# The Geometry of Concepts: Sparse Autoencoder Feature Structure

### v1

@@ Abstract
Sparse autoencoders have recently produced dictionaries of high-dimensional vectors corresponding to the universe of concepts represented by large language models. We find that this concept universe has interesting structure at three levels: 1) The “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen). We find that the quality of such parallelograms and associated function vectors improves greatly when projecting out global distractor directions such as word length, which is efficiently done with linear discriminant analysis. 2) The “brain” intermediate-scale structure has significant spatial modularity; for example, math and code features form a “lobe” akin to functional lobes seen in neural fMRI images. We quantify the spatial locality of these lobes with multiple metrics and find that clusters of co-occurring features, at coarse enough scale, also cluster together spatially far more than one would expect if feature geometry were random. 3) The “galaxy” scale large-scale structure of the feature point cloud is not isotropic, but instead has a power law of eigenvalues with steepest slope in middle layers. We also quantify how the clustering entropy depends on the layer.
@@ 1 Introduction
The past year has seen a breakthrough in understanding how large language models work: sparse autoencoders have discovered large numbers of points (“features”) in their activation space that can be interpreted as concepts (Huben et al., 2023; Bricken et al., 2023). Such SAE point clouds have recently been made publicly available (Lieberum et al., 2024), so it is timely to study their structure at various scales. [^1] Throughout the paper, we use the JumpReLU SAEs (Rajamanoharan et al., 2024) of “Gemma Scope” (Lieberum et al., 2024) for gemma-2-2b and gemma-2-9b (Team et al., 2024). By “SAE point cloud”, we refer to the collection of unit-norm decoder feature vectors learned by the SAE. Unless otherwise stated, we use residual stream SAEs ( 2b-pt-res or 9b-pt-res) with 16k features. This is the goal of the present paper, focusing on three separate spatial scales. In Section 3, we investigate if the “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen). In Section 4, we test if the “brain” intermediate-scale structure has functional modularity akin to biological brains. In Section 5, we study the “galaxy” scale large-scale structure of the feature point cloud, testing whether it is more interestingly shaped and clustered than an isosropic Gaussian distribution.
@@ 2 Related work
SAE feature structure: Sparse autoencoders have relatively recently garned attention as an approach for discovering interpretable language model features without supervision, with relatively few works examining SAE feature structure. Bricken et al. (2023) and Templeton et al. (2024) both visualized SAE features with UMAP projections and noticed that features tend to group together in “neighborhoods” of related features, in contrast to the approximately-orthogonal geometry observed in the toy model of Elhage et al. (2022). Engels et al. (2024) find examples of SAE structure where multiple SAE features appear to reconstruct a multi-dimensional feature with interesting geometry, and multiple authors have recently speculated that SAE vectors might contain more important structures (Mendel, 2024; Smith, 2024). Bussmann et al. (2024) suggest that SAE features are in fact linear combinations of more atomic features, and discover these more atomic latents with “meta SAEs”. Our discussion of crystal structure in SAE features is related to this idea.
@@
Function vectors and Word embedding models: Early word embedding methods such as GloVe and Word2vec, were found to contain directions encoding semantic concepts, e.g. the well-known formula f(king) - f(man) + f(woman) = f(queen) (Drozd et al., 2016; Pennington et al., 2014; Ma & Zhang, 2015). More recent research has found similar evidence of linear representations in sequence models trained only on next token prediction, including Othello board positions (Nanda et al., 2023; Li et al., 2022), the truth value of assertions (Marks & Tegmark, 2023), and numeric quantities such as longitude, latitude, birth year, and death year (Gurnee & Tegmark, 2023; Heinzerling & Inui, 2024). Recent works have found causal function vectors for in-context learning (Todd et al., 2023; Hendel et al., 2023; Kharlapenko et al., 2024), which induce the model to perform a certain task. Our discussion of crystal structures builds upon these previous works of finding task vectors or parallelogram structures in language models.
3 “Atom” scale: crystal structure
Figure 1: Parallelogram and trapezoid structure is revealed (left) when using LDA to project out distractor dimensions, tightening up clusters of pairwise Gemma-2-2b activation differences (right).
@@
In this section, we search for what we term crystal structure in the point cloud of SAE features. By this we mean geometric structure reflecting semantic relations between concepts, generalizing the classic example of $({\bf a},{\bf b},{\bf c},{\bf d})$=(man,woman,king,queen) forming an approximate parallelogram where ${\bf b}-{\bf a}\approx{\bf d}-{\bf c}$. This can be interpreted in terms of two function vectors ${\bf b}-{\bf a}$ and ${\bf c}-{\bf a}$ that turn male entities female and turn entities royal, respectively. We also search for trapezoids with only one pair of parallel edges ${\bf b}-{\bf a}\propto{\bf d}-{\bf c}$ (corresponding to only one function vector); Fig. 1 (right) shows such an example with $({\bf a},{\bf b},{\bf c},{\bf d})$=(Austria,Vienna,Switzerland,Bern), where the function vector can be interpreted as mapping countries to their capitals.
@@
We search for crystals by computing all pairwise difference vectors and clustering them, which should result in a cluster corresponding to each function vector. Any pair of difference vectors in a cluster should form a trapezoid or parallelogram, depending on whether the difference vectors are normalized or not before clustering (or, equivalently, whether we quantify similarity between two difference vectors via Euclidean distance or cosine similarity).
@@
Our initial search for SAE crystals found mostly noise. To investigate why, we focused our attention on Layers 0 (the token embedding) and 1, where many SAE features correspond to single words. We then studied Gemma2-2b residual stream activations for previously reported word $\mapsto$ word function vectors from the dataset of (Todd et al., 2023), which clarified the problem. Figure 1 illustrates that candidate crystal quadruplets are typically far from being parallelograms or trapezoids. This is consistent with multiple papers pointing out that (man,woman,king,queen) is not an accurate parallelogram either.
@@
We found the reason to be the presence of what we term distractor features. For example, we find that the horizontal axis in Figure 1 (right) corresponds mainly to word length (Appendix B, Figure 10), which is semantically irrelevant and wreaks havoc on the trapezoid (left), since “Switzerland” is much longer than the other words.
@@
To eliminate such semantically irrelevant distractor vectors, we wish to project the data onto a lower-dimensional subspace orthogonal to them. For the (Todd et al., 2023) dataset, we do this with Linear Discriminant Analysis (LDA) (Xanthopoulos et al., 2013), which projects onto signal-to-noise eigenmodes where “signal” and “noise” are defined as the covariance matrices of inter-cluster variation and intra-cluster variation, respectively. Figure 1 illustrates that this dramatically improves the cluster and trapezoid/parallelogram quality, highlighting that distractor features can hide existing crystals.
4 “Brain” scale: Meso-scale modular structure
Figure 2: Features in the SAE point cloud identified that tend to fire together within documents are seen to also be geometrically co-located in functional “lobes”, here down-projected to 2D with t-SNE with point size proportional to feature frequency. A 2-lobe partition (left) is seen to break the point cloud into roughly equal parts, active on code/math documents and English language documents, respectively. A 3-lobe partition (right) is seen to mainly subdivide the English lobe into a part for short messages and dialogue (e.g. chat rooms and parliament proceedings) and one primarily containing long-form scientific papers.
We now zoom out and look for larger-scale structure. In particular, we investigate if functionally similar groups of SAE features (which tend to fire together) are also geometrically similar, forming “lobes” in the activation space.
In animal brains, such functional groups are well-known clusters in the 3D space where neurons are located. For example, Broca’s area is involved in speech production, the auditory cortex processes sound, and the amygdala is primarily associated with processing emotions. We are curious whether we can find analogous functional modularity in the SAE feature space.
@@
We test a variety of methods for automatically discovering such functional “lobes” and for quantifying if they are spatially modular. We define a lobe partition as a partition of the point cloud into $k$ subsets (“lobes”) that are computed without positional information. Instead, we identify such lobes based on then being functionally related, specifically, tending to fire together within a document.
@@
To automatically identify functional lobes, we first compute a histogram of SAE feature co-occurrences. We take gemma-2-2b and pass documents from The Pile Gao et al. (2020) through it. In this section, we report results with a layer 12 residual stream SAE with 16k features and average L0 of 41. For this SAE, we record the features that fire (we count a feature as firing if its hidden activation $>1$). Features are counted as co-occurring if they both fire within the same block of 256 tokens – this length provides a coarse “time resolution” allowing us to find tokens that tend to fire together within the same document rather than just at the same token. We use a max context length of 1024, and only use one such context per document, giving us at most 4 blocks (and histogram updates) per document of The Pile. We compute histograms across 50k documents. Given this histogram, we compute an affinity score between each pair of SAE features based on their co-occurrence statistics and perform spectral clustering on the resulting affinity matrix.
@@
We experiment with the following notions of co-occurrence-based affinity: simple matching coefficient, Jaccard similarity, Dice coefficient, overlap coefficient, and Phi coefficient, which can all be computed just from a co-occurrence histogram. In the Appendix A.1, we review definitions for each of these and in Figure 5 illustrate how the choice between them affects the resulting lobes.
@@
Our null hypothesis is that functionally similar points (of commonly co-occurring SAE features) are uniformly distributed throughout the activation space, showing no spatial modularity. In contrast, Figure 2 shows lobes that appear visually quite spatially localized. To quantify how statistically significant this is, we use two approaches to rule out the null hypothesis.
@@
    While we can cluster features based on whether they co-occur, we can also perform spectral clustering based on the cosine similaity between SAE feature decoder vectors. Given a clustering of SAE features using cosine similarity and a clustering using co-occurrence, we compute the mutual information between these two sets of labels. In some sense, this directly measures the amount of information about geometric structure that one gets from knowing functional structure. We report the adjusted mutual information Vinh et al. (2009) as implemented by scikit-learn Pedregosa et al. (2011), which corrects for chance agreements between the clusters.
@@
    Another conceptually simple approach is to train models to predict which functional lobe a feature is in from its geometry. To do this, we take a given set of lobe labels from our co-occurrence-based clustering, and train a logistic regression model to predict these labels directly from the point positions, using an 80-20 train-test split and reporting the balanced test accuracy of this classifier.
Figure 3 shows that for both measures, the Phi coefficient wins, delivering the best correspondence between functional lobes and feature geometry. To show that this is statistically significant, we randomly permute the cluster labels from the cosine similarity-based clustering and measure the adjusted mutual information. We also randomly re-initialize the SAE feature decoder directions from a random Gaussian and normalize, and then train logistic regression models to predict functional lobe from these feature directions. Figure 3 (bottom) shows that both tests rule out the null hypothesis at high significance, at 954 and 74 standard deviations, respectively, clearly demonstrating that the lobes we see are real and not a statistical fluke.
To assess what each lobe specializes in, we run 10k documents from The Pile through gemma-2-2b, and again record which SAE features at layer 12 fire within blocks of 256 tokens. For each block of tokens, we record which lobe has the highest proportion of its features firing. Each document in The Pile is attached with a name specifying the subset of the corpus that document is from. For each document type, for each 256-token block within a document of that type, we record which lobe which had the highest proportion of its SAE features firing. Across thousands of documents, we can then look at a histogram of which lobes were maximally activating across each document type. We show these results for three lobes, computed with the Phi coefficient as the co-occurrence measure, in Figure 4. This forms the basis for our lobe labeling in Figure 2.
The effects of the five different co-occurrence measures are compared in Figure 5. Although we found Phi to be best, all five are seen to discover the “code/math lobe”.
Figure 3: Top Left: Adjusted mutual information between spatial clusters and functional (cooccurrence-based) clusters. Top Right: logistic regression balanced test accuracy, predicting cooccurrence-based cluster label from position. Bottom Left: Adjusted mutual information with randomly-permuted cosine similarity-based clustering labels. Bottom Right: balanced test accuracy with random unit-norm feature vectors. The statistical significance reported is for Phi-based clustering with into lobes.
@@
Figure 5: Comparison of the lobe partitions of the SAE point cloud discovered with different affinity measures, with the same t-SNE projection as Figure 2. In the top left, we show clusters computed from geometry, the angular similarity between features as the affinity score for spectral clustering. All other measures are based on whether SAE features co-occur (fire together) within 256-token blocks, using different measures of affinity. Although Phi predicts spatial structure best, all co-occurrence measures are seen to discover the “code/math lobe”.
5 “Galaxy” scale: Large-scale point cloud structure
In this section, we zoom out further and study the ”galaxy” scale structure of the point cloud, mainly its overall shape and clustering, analogously to how astronomers study galaxy shapes and substructure.
The simple null hypothesis that we try to rule out is that the point cloud is simply drawn from an isotropic multivariate Gaussian distribution. Figure 6 visually suggests that the cloud is not quite round even in its three first principal components, with some principal axes slightly wider than others akin to a human brain.
Figure 6: 3D Point Cloud visualizations of top PCA components for the Gemma2-2b layer 12 SAE features.

### v2

# sparse coding; mechanistic interpretability; neural networks; large language models; clustering

@@ Abstract
Sparse autoencoders have recently produced dictionaries of high-dimensional vectors corresponding to the universe of concepts represented by large language models. We find that this concept universe has interesting structure at three levels: (1) The “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen). We find that the quality of such parallelograms and associated function vectors improves greatly when projecting out global distractor directions such as word length, which is efficiently performed with linear discriminant analysis. (2) The “brain” intermediate-scale structure has significant spatial modularity; for example, math and code features form a “lobe” akin to functional lobes seen in neural fMRI images. We quantify the spatial locality of these lobes with multiple metrics and find that clusters of co-occurring features, at coarse enough scale, also cluster together spatially far more than one would expect if feature geometry were random. (3) The “galaxy”-scale large-scale structure of the feature point cloud is not isotropic, but instead has a power law of eigenvalues with steepest slope in middle layers. We also quantify how the clustering entropy depends on the layer.
keywords:
sparse coding; mechanistic interpretability; neural networks; large language models; clustering
[…]
@@ 1 Introduction
While large language models (LLMs) now exhibit a variety of impressive abilities Hurst et al. (2024); The ; Guo et al. (2025), we largely do not understand the internal cognition that underlies the behavior of these systems. This lack of transparency may pose a challenge for a variety of AI safety Slattery et al. (2024) concerns. For instance, it may be difficult to tell whether seemingly benign model behavior in any particular instance is sycophantic Sharma et al. (2023) or deceptive Park et al. (2023) without an analysis of the internals of the system. Such “interpretability” analysis has already shown promise in auditing AI systems Marks et al. (2024) to identify misaligned goals Ngo et al. (2022). As systems become more powerful, there is a need for methods to further our understanding of the internal representations and algorithms learned by these systems Bereska and Gavves (2024); Sharkey et al. (2025).
The past year has seen a breakthrough in understanding how large language models work: sparse autoencoders (SAEs) have discovered large numbers of vectors (“features”) in their activation space that can be interpreted as concepts (Huben et al., 2023; Bricken et al., 2023; Templeton et al., 2024). These advances build on earlier studies applying sparse coding to artificial neural network representations Faruqui et al. (2015); Zhang et al. (2019); Yun et al. (2021), and to earlier work in neuroscience on biological neural representations Olshausen and Field (1996, 1997). Underlying this work is the idea that neural networks use sparse coding to represent concepts in their activation space Elhage et al. (2022). In particular, sparse autoencoders are motivated by the assumptions that (1) networks compute a variety of “features” from their input, (2) features are represented as one-dimensional directions in activation space $\{\bm{d}_{i}\}$, (3) features are represented simply by adding them to the network’s activations, so activation vectors take the form $\sum_{i}f_{i}\bm{d}_{i}$, and (4) the coefficients $f_{i}$ are sparse—only a small subset of all possible features “fire” at once. The combination of assumptions (2)–(4) has been called the Linear Representation Hypothesis Park et al. (2023); Olah (2024); Engels et al. (2024).
If these assumptions hold, we could automatically discover these features with sparse dictionary learning. Sparse dictionary learning attempts to learn an overcomplete basis (dictionary) $\{\bm{d}_{i}\}$ such that vectors $\bm{x}$ from a given distribution can be represented as sparse linear combinations of dictionary elements. Sparse autoencoders offer a simple approach to sparse dictionary learning. Sparse autoencoders consist of a learnable encoder function Enc, which maps vectors $\bm{x}\in\mathbb{R}^{n}$ to a hidden latent representation $\bm{f}\in\mathbb{R}^{m}$, and a decoder Dec, which maps latent $\bm{f}$ back to $\bm{\hat{x}}\in\mathbb{R}^{n}$. The objective of the sparse autoencoder is to accurately reconstruct the input $\bm{x}$ from a sparse latent representation, and they are trained with gradient descent with a loss function like
$\mathcal{L}=||\bm{x}-\text{Dec}(\text{Enc}(\bm{x}))||_{2}^{2}+\lambda||\bm{f}||_{0}.$
Sparse autoencoders use a linear decoder $\text{Dec}(\bm{f})=\bm{W}_{d}\bm{f}+b_{d}$, so that the output of the SAE can be interpreted as a linear combination of features: $\bm{\hat{x}}=\sum_{i}f_{i}\bm{W}_{d}^{i}+\bm{b}_{d}$. In practice, hidden latents discovered by sparse autoencoders tend to be more interpretable than neurons, activating in more consistent contexts Huben et al. (2023); Bricken et al. (2023), suggesting that they may be learning the true latents underlying the network’s computation. For AI safety, sparse autoencoders have shown some preliminary success: Ref. Marks et al. (2024) reports specially training an LLM to have a hidden objective, and then challenging separate teams of researchers to identify this objective. One team was able to quickly identify this objective by looking at sparse autoencoder features that activated when the LLM was prompted to exhibit “potentially concerning behaviors”, and then looking at examples in the training data where that same feature fired.
Although some early work motivating sparse autoencoders suggested that networks would arrange features maximally spread apart (approximately orthogonal) Elhage et al. (2022), recent works have suggested that features may have a more sophisticated geometric structure Engels et al. (2024); Templeton et al. (2024). Recently, a large collection of SAEs have been made publicly available Lieberum et al. (2024), so it is timely to study their structure at various scales. Thus, the present paper examines sparse autoencoder feature structure at three separate spatial scales, which we refer to informally as the “atom”-scale, “brain”-scale, and “galaxy”-scale. These playful analogies are not meant to be precise, but instead gesture at certain concepts and methods of analysis from other fields which we apply to understanding language model feature structure. We provide project code at https://github.com/ejmichaud/feature-geometry (accessed on 24 March 2025).
@@
This paper is organized as follows. In Section 2, we summarize related work. In Section 3, we investigate if the “atomic” small-scale structure contains “crystals” whose faces are parallelograms or trapezoids, generalizing well-known examples such as (man:woman::king:queen). In Section 4, we test if the “brain” intermediate-scale structure has functional modularity akin to biological brains. In Section 5, we study the “galaxy” large-scale structure of the feature point cloud, testing whether it is more interestingly shaped and clustered than an isotropic Gaussian distribution, and conclude in Section 6.
@@ 2 Related Work
Neural network geometry: Many past works have studied the geometry of neural network activations. These works find that the intrinsic dimension of neural network hidden states are much lower than the full model dimension Ansuini et al. (2019), that nearby vectors in activation space are semantically similar Chandrasekaran and Mago (2021), and that at local minima well generalizing neural network loss landscapes have many “flat” directions Watanabe (2009). Other works study how representations evolve through models; one hypotheses is “iterative inference”, which claims that neural networks iteratively refine activations layer by layer Rushing and Nanda (2024); Belrose et al. (2023). A contrasting hypothesis is a circuits view, which holds that information flows in discrete steps along a directed acyclic graph through the model, and representations cleanly change between steps Conmy et al. (2023). Another work Park et al. (2024) found that representations of hierarchically related concepts are orthogonal to each other while categorical concepts are represented as polytopes. Our work is in the same vein as these earlier analysis, but differs in an important way because we use the SAE basis, which represents the model’s atomic concept space instead of its activation space.
@@
SAE feature structure: Sparse autoencoders (SAEs) are a recent approach for discovering interpretable language model features without supervision, although relatively few works have examined SAE feature structure. Bricken et al. (2023) and Templeton et al. (2024) both visualize SAE features with UMAP projections and notice that features tend to group together in “neighborhoods” of related features, in contrast to the approximately orthogonal geometry observed in the toy model of Elhage et al. (2022). Engels et al. (2024) find examples of SAE structure where multiple SAE features appear to reconstruct a multi-dimensional feature with interesting geometry, and multiple authors have recently speculated that SAE vectors might contain more important structures (Mendel, 2024; Smith, 2024). Bussmann et al. (2024) suggest that SAE features are in fact linear combinations of more atomic features, and discover these more atomic latents with “meta SAEs”. Our discussion of crystal structure in SAE features is related to this idea that seemingly atomic representations might be composed of more atomic representations.
@@
Semantically meaningful linear representations: Early work found that word embedding methods such as GloVe and Word2vec contained directions encoding semantic concepts, e.g., the well-known formula f(king) -- f(man) + f(woman) = f(queen) (Drozd et al., 2016; Pennington et al., 2014; Ma and Zhang, 2015). More recent research has found similar evidence of linear representations in sequence models trained only on next token prediction, including Othello board positions (Nanda et al., 2023; Li et al., 2022), integer lattices (Michaud et al., 2024), the truth value of assertions (Marks and Tegmark, 2023), and numeric quantities such as longitude, latitude, birth year, and death year (Gurnee and Tegmark, 2023; Heinzerling and Inui, 2024), inspiring the Linear Representation Hypothesis (see above). Recent works have also found causal function vectors for in-context learning (Todd et al., 2023; Hendel et al., 2023; Kharlapenko et al., 2024). These function vectors induce the model to perform a certain task when added into the model’s hidden states. Our discussion of crystal structures builds upon these previous works by finding these task vectors and parallelogram structures in sparse autoencoder representations.
3 “Atom”-Scale: Crystal Structure
@@
In this section, we search for what we term crystal structure in the point cloud of SAE features. By this we mean geometric structure reflecting semantic relations between concepts, generalizing the classic example of $({\bf a},{\bf b},{\bf c},{\bf d})$=(man,woman,king,queen) forming an approximate parallelogram where ${\bf b}-{\bf a}\approx{\bf d}-{\bf c}$. This can be interpreted in terms of two function vectors ${\bf b}-{\bf a}$ and ${\bf c}-{\bf a}$ that turn male entities female and turn entities royal, respectively. We also search for trapezoids with only one pair of parallel edges ${\bf b}-{\bf a}\propto{\bf d}-{\bf c}$ (corresponding to only one function vector); Figure 1 (right) shows such an example with $({\bf a},{\bf b},{\bf c},{\bf d})$=(Austria,Vienna,Switzerland,Bern), where the function vector can be interpreted as mapping countries to their capitals. Studying these crystal structures is important because they provide insight into how LLMs internally represent semantic operations and relational knowledge. For instance, function vectors help us assess the extent to which semantic structures within models align with human intuition and language logic, as explored in the recent literature Todd et al. (2023); Meng et al. (2022).
Figure 1: Parallelogram and trapezoid structure is revealed (left) when distractor dimensions were projected out from the activations using LDA. LDA results in tighter clusters of pairwise Gemma-2-2b activation differences (right), where each cluster corresponds to a different semantic transformation. Distractor features are defined as those that are not related to semantics of the text; for instance, the first principal component of Gemma-2-2b’s Layer 0 activations (top left figure on the right panel) represents word length. Parallelogram or trapezoid structures suggest that there is a unique direction in the activation space that represents each semantic transformation.
@@
We search for crystals by computing all pairwise difference vectors and clustering them using the K-means algorithm Ahmed et al. (2020), where the vectors could be either the original model’s hidden state activations (model crystal) or SAE features’ decoder vectors (SAE crystal). We use Gemma-2-2b for the experiment. If there is a direction that represents each semantic transformation, we expect each resulting cluster to correspond to each function vector. In other words, any pair of difference vectors in a cluster will form a trapezoid or parallelogram, depending on whether the difference vectors are normalized or not before clustering (or, equivalently, whether we quantify similarity between two difference vectors via Euclidean distance or cosine similarity).
@@
Our initial search for SAE crystals found mostly noise. To investigate why, we decided to focus on activations of the model’s early layer, where many SAE features correspond to a single token. Since SAE feature vectors in the early layers are often closely related to the corresponding model activations, we believed that studying the activations of these early layers could help clarify why our initial crystal search primarily found noise. Therefore, we studied Gemma-2-2b residual stream activations for previously reported word $\mapsto$ word function vectors from the dataset of (Todd et al., 2023). Figure 1 illustrates that candidate crystal quadruplets are typically far from being parallelograms or trapezoids. This is consistent with multiple papers pointing out that (man,woman,king,queen) is not an accurate parallelogram either.
@@
We believe the reason to be the presence of what we term distractor features. We define distractor features to be the features that are not related to semantics of the text. For example, we find that the horizontal axis in Figure 1 (right) corresponds mainly to word length (Appendix B, Figure 10), which is semantically irrelevant and wreaks havoc on the trapezoid (left), since “Switzerland” is much longer than the other words. However, these distractor features were not always interpretable; in some cases, it was difficult to associate features with any clear linguistic property.
@@
To eliminate such semantically irrelevant distractor vectors, we wish to project the data onto a lower-dimensional subspace orthogonal to them. For the (Todd et al., 2023) dataset, we accomplish this with linear discriminant analysis (LDA) (Xanthopoulos et al., 2013), which projects onto signal-to-noise eigenmodes where “signal” and “noise” are defined as the covariance matrices of inter-cluster variation and intra-cluster variation, respectively. Figure 1 illustrates that this dramatically improves the cluster and trapezoid/parallelogram quality, highlighting that distractor features can hide existing crystals.
4 “Brain”-Scale: Meso-Scale Modular Structure
We now zoom out and look for larger-scale structure. In particular, we investigate if functionally similar groups of SAE features (which tend to fire together) are also geometrically similar, forming “lobes” in the activation space. We refer to this analyis as “brain”-scale because, in animal brains, functionally similar groups of neurons also typically cluster together spatially. For example, Broca’s area is involved in speech production, the auditory cortex processes sound, and the amygdala is primarily associated with processing emotions. We are curious whether we can find analogous functional modularity in the SAE feature space. While prior work has qualitatively observed that semantically related features are spatially close via UMAP projections of features Bricken et al. (2023); Templeton et al. (2024), we aim to more precisely quantify the relationship between functional similarity and spatial similarity.
@@
We test a variety of methods for automatically discovering such functional “lobes” and for quantifying if they are spatially modular. We define a lobe partition as a partition of the SAE feature point cloud into $k$ subsets (“lobes”) that are computed without positional information. Instead, we identify such lobes based on them being functionally related, specifically, tending to fire together within a document.
@@
To automatically identify functional lobes, we first compute a histogram of SAE feature co-occurrences. We take Gemma-2-2b and pass documents from The Pile Gao et al. (2020) through it. In this section, we report results with a Layer 12 residual stream SAE with 16k features and an average L0 of 41. For this SAE, we record the features that fire (we count a feature $i$ as firing if its encoder assigns it a coefficient $f_{i}>1$). Features are counted as co-occurring if they both fire within the same block of 256 tokens—this length provides a coarse “time resolution” allowing us to find tokens that tend to fire together within the same document rather than just at the same token. We use a max context length of 1024, and only use one such context per document, giving us at most four blocks (and histogram updates) per document of The Pile. We compute histograms across 50k documents. Given this histogram, we compute an affinity score between each pair of SAE features based on their co-occurrence statistics and perform spectral clustering on the resulting affinity matrix. We use the spectral clustering implementation of scikit-learn Pedregosa et al. (2011) with default settings with varying choice of n_clusters.
In Figure 2, we visualize lobes discovered with this method with n_clusters=2, 3 via a t-SNE projection Van der Maaten and Hinton (2008). For this figure, we used the “phi coefficent” as the measure of co-occurrence similarity between features. We find that lobes visually appear to be spatially localized. For instance, features which fire primarily on math and code documents tend to cluster together spatially.
@@
We experiment with the following notions of co-occurrence-based affinity: simple matching coefficient, Jaccard similarity, Dice coefficient, overlap coefficient, and phi coefficient, which can all be computed just from a co-occurrence histogram. In the Appendix A.1, we review definitions for each of these and in Figure 3 illustrate how the choice between them affects the resulting lobe t-SNE plots. We also show how lobes appear when we cluster based on geometry directly using cosine similarities, as described below.
Figure 2: Features in the SAE feature point cloud identified that tend to fire together within documents are seen to also be geometrically co-located in functional “lobes”, here down-projected to 2D with t-SNE with point size proportional to feature frequency. A 2-lobe partition (left) is seen to break the point cloud into roughly equal parts, active on code/math documents and English language documents, respectively. A 3-lobe partition (right) is seen to mainly subdivide the English lobe into a part for short messages and dialogue (e.g., chat rooms and parliament proceedings) and one primarily containing long-form scientific papers.
Figure 3: Comparison of the lobe partitions of the SAE point cloud discovered with different affinity measures, with the same t-SNE projection as Figure 2. In the top left, we show clusters computed from geometry, the cosine similarity between features as the affinity score for spectral clustering. All other measures are based on whether SAE features co-occur (fire together) within 256-token blocks, using different measures of affinity. Although the phi coefficient predicts spatial structure best, all co-occurrence measures are seen to discover the code/math lobe.
@@
While these plots show a qualitative relationship between co-occurrence and feature geometry, we aim to quantify this relationship. Our null hypothesis is that functionally similar points (of commonly co-occurring SAE features) are uniformly distributed throughout the activation space, showing no spatial modularity. To quantify how statistically significant this is, we use two approaches to rule out the null hypothesis:
@@
    While we can cluster features based on whether they co-occur, we can also perform spectral clustering based on the cosine similarity between SAE feature decoder vectors. So instead of feature affinity values being, e.g., their co-occurrence phi coefficient, affinity matrix values are instead computed simply from feature geometry as $A_{ij}=\bm{d}_{i}\cdot\bm{d}_{j}$. Given a clustering of SAE features using cosine similarity and a clustering using co-occurrence, we compute the mutual information between these two sets of labels. In some sense, this measures the amount of information about geometric structure that one obtains from knowing functional structure. We report the adjusted mutual information Vinh et al. (2009) as implemented by scikit-learn Pedregosa et al. (2011), which corrects for chance agreements between the clusters.
@@
    Another conceptually simple approach is to train models to predict which functional lobe a feature is in from its geometry. To accomplish this, we take a given set of lobe labels from our co-occurrence-based clustering, and train a logistic regression model to predict these labels directly from the point positions, using an 80-20 train–test split and reporting the balanced test accuracy of this classifier.
Figure 4 shows that for both measures, the phi coefficient gives the best correspondence between functional lobes and feature geometry. To show that this is statistically significant, we randomly permute the cluster labels from the cosine similarity-based clustering and measure the adjusted mutual information. We also re-initialize the SAE feature decoder directions from a random Gaussian and normalize, and then train logistic regression models to predict functional lobe from these random feature directions. Figure 4 (bottom) shows that both tests rule out the null hypothesis with high significance, at 954 and 74 standard deviations, respectively, clearly demonstrating that the lobes we see are real and not a statistical fluke.
Figure 4: (top left): Adjusted mutual information between spatial clusters and functional (co-occurrence-based) clusters. (top right): logistic regression balanced test accuracy, predicting co-occurrence-based cluster label from position. (bottom left): Adjusted mutual information with randomly permuted cosine similarity-based clustering labels. (bottom right): balanced test accuracy with random unit-norm feature vectors. The statistical significance reported is for phi-based clustering into lobes.
To assess what each lobe specializes in, we run 10k documents from The Pile through Gemma-2-2b, and again record which SAE features at Layer 12 fire within blocks of 256 tokens. For each block of tokens, we record which lobe has the highest proportion of its features firing. Each document in The Pile is attached with a name specifying the subset of the corpus that document is from. For each document type, for each 256-token block within a document of that type, we record which lobe had the highest proportion of its SAE features firing. Across thousands of documents, we can then look at a histogram of which lobes were maximally activating across each document type. We show these results for three lobes, computed with the phi coefficient as the co-occurrence measure, in Figure 5. This forms the basis for our lobe labeling in Figure 2.
These findings raise interesting questions about whether individual sparse autoencoder features are the most natural units for understanding neural networks Mueller et al. (2024); Olah (2023). In biological brains, one can study individual neurons, groups of neurons, groups of groups of neurons, and so on up to very large-scale structures, and it is not clear a priori what “scale” of analysis will be most fruitful Hoel et al. (2013). We may face a similar ambiguity with sparse autoencoder features, since, as we have seen, groups of co-occurring, geometrically related features can be interpretable and studied in their own right. This question, of whether there is a right “scale” of analysis for SAE features, is made even more salient by the observation in prior work of “feature splitting” Bricken et al. (2023).
@@
5 “Galaxy”-Scale: Large-Scale Point Cloud Structure
In this section, we further broaden our perspective to analyze the “galaxy”-scale structure of the point cloud, focusing on its overall shape and clustering properties. This analysis is loosely inspired by work in astronomy Kennicutt Jr (1998) characterizing the shape Hubble (1926) and substructure Kravtsov (2010) of galaxies. We start by formulating a simple null hypothesis: The point cloud is drawn from an isotropic multivariate Gaussian distribution.
To test this, we analyze the covariance of the data. As illustrated in Figure 6, the eigenvalue spectrum deviates from isotropy, meaning the cloud exhibits directional structure rather than being purely spherical. Even within the first three principlal components, the point cloud is anisotrophic, with some principal axes slightly wider than others.
To quantify these deviations, we analyze the eigenvalue spectrum of the covariance matrix, comparing it to theoretical expectations from random matrix theory (RMT).
Figure 6: Three-dimensional point cloud visualizations of top PCA components for the Gemma-2-2b Layer 12 SAE features.